# AED-PADA:Improving Generalizability of Adversarial Example Detection via Principal Adversarial Domain Adaptation

### v1

@@ Abstract
Adversarial example detection, which can be conveniently applied in many scenarios, is important in the area of adversarial defense. Unfortunately, existing detection methods suffer from poor generalization performance, because their training process usually relies on the examples generated from a single known adversarial attack and there exists a large discrepancy between the training and unseen testing adversarial examples. To address this issue, we propose a novel method, named Adversarial Example Detection via Principal Adversarial Domain Adaptation (AED-PADA). Specifically, our approach identifies the Principal Adversarial Domains (PADs), i.e., a combination of features of the adversarial examples from different attacks, which possesses large coverage of the entire adversarial feature space. Then, we pioneer to exploit multi-source domain adaptation in adversarial example detection with PADs as source domains. Experiments demonstrate the superior generalization ability of our proposed AED-PADA. Note that this superiority is particularly achieved in challenging scenarios characterized by employing the minimal magnitude constraint for the perturbations.
1 Introduction
@@
Recently, Deep Neural Networks (DNNs) have been playing prominent roles in many applications. Unfortunately, considerable studies have demonstrated that DNNs can be easily deceived if certain imperceptible perturbations are introduced to their inputs [?????]. These perturbed inputs, also known as adversarial examples, have enforced DNNs to produce erroneous decision and developed into a significant security concern in safety sensitive scenarios, such as autonomous driving [?] and medical diagnosis [?].
Figure 1: Schematic illustration of our proposed work. (a) represents the entire adversarial feature space as an instance, which contains 9 adversarial domains $\{A_{1},A_{2},\ldots,A_{9}\}$. (b) represents the mechanism of existing detection methods, which usually performs the training via a single source domain to detect the examples in the unseen adversarial domain, e.g. $T~{}(T=A_{9})$. (c) is a straightforward solution to improve generalization ability via randomly selecting multiple source domains. (d) presents the intuition of our work, which constructs PADs to possess a larger coverage of the entire feature space to create more potential overlaps with the target domain, which can better improve the generalization ability.
@@
Nowadays, adversarial training has been proved to be an effective adversarial defense strategy [???]. However, it requires to possess enough knowledge about the classification models and necessitates substantial computational costs to retrain the classification models. On the contrary, adversarial example detection, a.k.a. adversarial detection, defends the adversarial attacks by distinguishing whether the inputs are benign or manipulated. The mechanism of this type of methods can be efficiently deploy to defend many applications without the requirement of extra knowledge about the core model.
@@
Generalization ability is vital for adversarial detection methods in real-world scenarios, because these methods tend to encounter unseen attacks and they are desired to perform consistently. Current detection techniques [?????] typically achieve considerable generalization ability over a few conventional attacks, such as BIM [?], PGD [?] and C&W [?]. However, we empirically observe that these methods exhibit instability and inadequate performance against recent attacks, such as SSA [?], Jitter [?] and ILA-DA [?]. These methods tend to give unsatisfactory generalization performance, because their training process usually relies on a single known adversarial attack and they have not been developed from the perspective of boosting generalization ability.
@@
In this paper, we provide a new perspective to further analyze the generalization ability of adversarial detection methods. Specifically, we assume that all the examples generated from adversarial attacks share a high dimensional manifold, and the features of the examples generated from each attack form a distinct domain, named Adversarial Domain (AD), as depicted in Fig. 1(a). Formally, Adversarial Domain (AD) of a particular adversarial attack is defined as the cumulative representations of all the adversarial examples, which are generated by that attack. Existing detection methods typically select a random single attack to generate the training samples, i.e., they only select one AD as the source domain, as shown in Fig. 1(b). Apparently, there is few intersections between the source and unseen target domains, which usually induce poor generalization performance. Besides, the randomness inherent in the selection of source domain will induce considerable fluctuations in generalization performance.
@@
A straightforward solution to the above issues, as shown in Fig. 1(c), is to randomly select multiple ADs as the source domains. This strategy tends to create a larger overlap with the target domain(s) and thereby improves the generalization performance of adversarial detection. Nonetheless, this strategy also induces uncertainty, and the selected similar ADs bring additional training costs without enough performance gains.
@@
To further address the above problems, we propose a novel detection method, named Adversarial Example Detection via Principal Adversarial Domain Adaptation (AED-PADA). As shown in Fig. 1(d), by selecting Principal Adversarial Domains (PADs) as the source domains, which can significantly enlarges the coverage of the entire adversarial feature space and creates larger overlap with the target domain, we can offer superior generalization performance.
@@
Specifically, AED-PADA contains two stages, i.e., Principal Adversarial Domains Identification (PADI) and Principal Adversarial Domain Adaptation (PADA). In the stage of PADI, since the discrepancies between the adversarial examples from various adversarial attacks are quite different compared to these of the ordinary classification tasks, we exploit adversarial supervised contrastive learning(Adv-SCL) to construct distinguishable ADs. Then, we construct AD clustering to identify similar ADs. To select the most representative ADs of each cluster to form PADs, we propose a Coverage of Entire Feature Space (CEFS) metric. With our CEFS metric, the formed PADs possess a board coverage of the entire feature space, and thus improve the likelihood of capturing the location of unknown target AD(s). In the stage of PADA, to effectively utilize the rich knowledge acquired by PADs to detect the unseen adversarial examples in the target domain, we propose PADA, which is compatible with various existing Multi-source Domain Adaptation (MDA) methods, by pioneering to exploit the mechanism of MDA for adversarial example detection. Since typical MDAs only focus on extracting the semantic features from the spatial domain, we propose an adversarial feature enhancement module to extracts features from both the spatial and frequency domains to construct a more comprehensive representation of adversarial examples.
@@
  • We propose Principal Adversarial Domain Adaptation for detecting adversarial examples, by exploiting adversarial feature enhancement based Multi-source Domain Adaptation (MDA), which is compatible with various existing MDA methods. To the best of our knowledge, this is the first work which exploits MDA for adversarial example detection.
2 Related Work
Adversarial example detection.
@@ 2 Related Work
The majority of existing adversarial example detection methods rely on statistical features [??????]. They usually assume that the benign and adversarial examples originate from different distributions, and construct detectors based on the distinct statistical characteristics of these examples. Specifically, Grosse et al. [?] utilize Maximum Mean Discrepancy for the adversarial example detection. Feinman et al. [?] carry out the detection based on Kernel Density (KD) and Bayesian-Uncertainty (BU) estimation. Ma et al. [?] exploit the concept of Local Intrinsic Dimensionality (LID) to calculate the distance between the distribution of inputs and their neighbors. Lee et al. [?] utilize Guassian Discriminant Analysis (GDA) to model the difference between the benign and adversarial samples, and differentiate them based on Mahalanobis Distance (MD). Liu et al. [?] propose a steganalysis based detection method (Steg). Tian et al. [?] reveal the inconsistency in the boundary fluctuations between the adversarial and benign examples, and construct Sensitivity Inconsistency Detector (SID) to identify the adversarial examples. Wang et al. [?] embed hidden-layer feature maps of DNNs into word vectors, and detecting adversarial examples via Sentiment Analysis (SA).
Figure 2: Our AED-PADA framework contains two stages: (a) Principal Adversarial Domains Identification, which consists of Adversarial Domain Acquisition, Adversarial Domain Clustering and Principal Adversarial Domains Selection, and (b) Principal Adversarial Domain Adaptation.
3 Methodology
@@
To improve the generalization ability of adversarial example detection, we propose a novel adversarial example detection method, named Adversarial Example Detection via Principal Adversarial Domain Adaptation (AED-PADA). The entire framework of our AED-PADA is shown in Fig. 2. AED-PADA contains two stages, Principal Adversarial Domains Identification (PADI) and Principal Adversarial Domain Adaptation (PADA). In the stage of PADI, we firstly incorporate adversarial supervised contrastive learning(Adv-SCL) to acquire distinguishable ADs. Then, we construct AD clustering to group ADs into different clusters. By proposing the Coverage of Entire Feature Space (CEFS) metric, we select the most representative ADs from each cluster to form PADs. In the stage of PADA, we propose an adversarial feature enhancement based MDA method to effectively leverage PADs to achieve adversarial detection. Note that Secs. 3.2, 3.3, and 3.4 introduce PADI while Sec. 3.5 presents PADA.
3.1 Notations
@@
Suppose we have a labeled $C-$class classification dataset $\mathbb{D}=\{(x,y)\}_{i=1}^{N}$ with $N$ samples, where the label $y\in\{1,2,\ldots,C\}$. A classifier $f$ is trained on $\mathbb{D}$ to classify an input sample into one of the $C$ classes, $f(x)\rightarrow\mathbb{Z}_{C}$. We have a set of adversarial attacks, $\mathbb{A}=\{Adv_{1},Adv_{2},\ldots,Adv_{M}\}$, which consists of $M$ distinct adversarial attack methods. Each of them aims to deceive $f$ into assigning incorrect labels via the generated adversarial examples. $\mathbb{D}_{m}=\{(x^{\prime},y^{\prime},\psi)\}_{i=1}^{N}$ is the dataset generated by attack $Adv_{m}$, where $x^{\prime}$ and $y^{\prime}$ denote the adversarial examples and their corresponding predictions, respectively, and $\psi$ presents the attack employed to generate $x^{\prime}$. $\psi=m$, for all $\psi$ in $\mathbb{D}_{m}$. Consequently, we construct a set of adversarial examples $\mathbb{O}$, which comprises $M$ types of adversarial examples, $\mathbb{O}=\{\mathbb{D}_{1},\mathbb{D}_{2},\ldots,\mathbb{D}_{M}\}$.
3.2 Adversarial Domain Acquisition
@@
Typically, if we extract the features of the adversarial examples, which are generated from different untargeted attacks, via common CNNs, the features tend to spread in an indistinguishable manner in the feature space. To acquire distinguishable representations of adversarial examples from different attacks, we exploit adversarial supervised contrastive learning (Adv-SCL) to extract features. Then, we form the Adversarial Domains (ADs), each of which is defined as the cumulative representations of all the adversarial examples generated from a particular adversarial attack.
@@
Based on the supervised contrastive learning (SCL) method [?], Adv-SCL neglects the classification result $y^{\prime}$ of adversarial examples and only focuses on identifying their generation methods. Specifically, each adversarial example in $\mathbb{O}$ is characterized by its generation method, denoted as $\psi$, which serves as a key discriminant between the positive and negative samples. Here, a pair of examples from identical attack are considered as positive, while those from different attacks are considered as negative. This learning strategy amplifies the dissimilarities across examples from various attacks and generates a more appropriate representation.
@@
As shown in Fig. 2(a), the input of AD acquisition is the adversarial example set $\mathbb{O}$. Adv-SCL consists of Encoder Network ($Enc$) and Projection Network ($Proj$). $Enc(\cdot)$ extracts a feature vector from the input image $x^{\prime}$, and $Proj(\cdot)$ further project this representation vector to an auxiliary vector $z=Proj(Enc(x^{\prime}))$, which will be discarded after training.
@@
For $n$ randomly sampled example-label pairs, $\{x_{k}^{\prime},y_{k}^{\prime},\psi_{k}\}_{k=1}^{n}$, in $\mathbb{O}$, the corresponding batch employed for training consists of $2n$ pairs, $\left\{\widetilde{x_{l}^{\prime}},\widetilde{y_{l}^{\prime}},\widetilde{\psi_{l}}\right\}_{l=1}^{2n}$, where $\widetilde{x^{\prime}}_{2k}$ and $\widetilde{x^{\prime}}_{2k-1}$ are two views of $x^{\prime}_{k}$, and $\widetilde{\psi}_{2k-1}=\widetilde{\psi}_{2k}=\psi_{k}$. The loss of Adv-SCL is defined as,
$\mathcal{L}^{sup}=\sum_{i\in I}\mathcal{L}_{i}^{sup},$ (1)
$\mathcal{L}_{i}^{sup}=\frac{-1}{|P(i)|}\sum_{p\in P(i)}\log\frac{\exp(z_{i}\cdot z_{p}/\tau)}{\sum_{a\in A(i)}\exp(z_{i}\cdot z_{a}/\tau)}.$ (2)
@@
Here, $i\in I=\{1,\ldots,2N\}$ denotes the index of the augmented samples. $A(i)$ is a subset of $I$ containing all the indexes except $i$. $P(i)\equiv\{p\in A(i):\widetilde{\psi}_{p}=\widetilde{\psi}_{i}\}$ represents the indices of positive samples in the batch except $i$, and $|P(i)|$ stands for its cardinality. $\cdot$ denotes the dot product. $\tau$ is the temperature parameter which scales the similarity values.
3.3 Adversarial Domain Clustering

### v2

@@ Abstract.
Adversarial example detection, which can be conveniently applied in many scenarios, is important in the area of adversarial defense. Unfortunately, existing detection methods suffer from poor generalization performance, because their training process usually relies on the examples generated from a single known adversarial attack and there exists a large discrepancy between the training and unseen testing adversarial examples. To address this issue, we propose a novel method, named Adversarial Example Detection via Principal Adversarial Domain Adaptation (AED-PADA). Specifically, our approach identifies the Principal Adversarial Domains (PADs), i.e., a combination of features of the adversarial examples generated by different attacks, which possesses a large portion of the entire adversarial feature space. Subsequently, we pioneer to exploit Multi-source Unsupervised Domain Adaptation in adversarial example detection, with PADs as the source domains. Experimental results demonstrate the superior generalization ability of our proposed AED-PADA. Note that this superiority is particularly achieved in challenging scenarios characterized by employing the minimal magnitude constraint for the perturbations.
Adversarial defense, adversarial example detection, generalization ability.
copyright: acmlicensedjournal: TOMMjournalyear: 2024doi: 10.1145/3706061isbn: 978-1-4503-XXXX-X/18/06ccs: Security and privacy Social aspects of security and privacyccs: Computing methodologies Computer vision
1. Introduction
@@
Recently, Deep Neural Networks (DNNs) have been playing prominent roles in many applications. Unfortunately, considerable studies have demonstrated that DNNs can be easily deceived if certain imperceptible perturbations are introduced to their inputs (SzegedyZSBEGF13, ; NguyenYC15, ; shi2021, ; SSA, ; ILA-DA, ; a2sc, ). These perturbed inputs, also known as adversarial examples, have enforced DNNs to produce erroneous decision and become a significant security concern in safety sensitive scenarios, such as autonomous driving (attack_driving, ) and medical diagnosis (attack_medical, ).
@@
Nowadays, adversarial training has been proved to be an effective adversarial defense strategy (FGSM, ; AD_tripletloss, ; XieTGWYL20, ; adversarial_training_tomm, ). However, it requires to possess enough knowledge about the classification models and necessitates substantial computational costs to retrain the classification models. On the contrary, adversarial example detection, a.k.a. adversarial detection, defends against adversarial attacks by distinguishing whether the inputs are benign or manipulated. The mechanism of this type of methods can be efficiently deploy to defend many applications without the requirement of extra knowledge about the core model.
@@
Generalization ability is vital for adversarial detection methods in real-world scenarios, because these methods tend to encounter unseen attacks and they are desired to perform consistently. Current detection techniques (lid, ; md, ; steg, ; SID, ; txt_advdetection, ) typically achieve considerable generalization ability over a few conventional attacks, such as BIM (BIM, ), PGD (pgd, ), and C&W (CW, ). However, we empirically observe that these methods exhibit instability and inadequate performance against recent attacks, such as SSA (SSA, ), Jitter (jitter, ) and ILA-DA (ILA-DA, ). These methods tend to give unsatisfactory generalization performance, because their training process usually relies on a single known adversarial attack and they have not been developed from the perspective of boosting generalization ability.
@@
In this paper, we provide a new perspective to further analyze the generalization ability of adversarial detection methods. Clearly, the threat model of all different adversarial attacks should be identical to ensure fair analysis and comparison. Thus, the features extracted by the model from different attacks are of same dimensions within a shared feature space. Due to the variations in configurations such as attack objectives, parameters, and loss functions, the features of the examples generated from each attack form a distinct domain, named Adversarial Domain (AD), as depicted in Fig. 1(a). Formally, Adversarial Domain (AD) of a particular adversarial attack is defined as the cumulative representations of all the adversarial examples, which are generated by that attack. Existing detection methods typically select a random single attack to generate the training samples, i.e., they only select one AD as the source domain, as shown in Fig. 1(b). Apparently, there are few intersections between the source and unseen target domains, which usually lead to poor generalization performance. Additionally, the randomness inherent in the selection of the source domains will induce considerable fluctuations in generalization performance.
Figure 1. Schematic illustration of our proposed work. (a) represents the entire adversarial feature space as an instance, which contains 9 adversarial domains $\{A_{1},A_{2},\ldots,A_{9}\}$. (b) represents the mechanism of existing detection methods, which usually performs the training via a single source domain to detect the examples in the unseen adversarial domain, e.g. $T~{}(T=A_{9})$. (c) is a straightforward solution to improve generalization ability via randomly selecting multiple source domains. (d) presents the intuition behind our work. We construct PADs, which possess a larger coverage of the entire feature space, to create more potential overlaps with the target domain. The strategy is designed to significantly enhance the detection generalization ability.
@@
A straightforward solution to the above issues, as shown in Fig. 1(c), is to randomly select multiple ADs as the source domains. This strategy tends to create a larger overlap with the target domain(s) and thereby improves the generalization performance of adversarial detection. Nonetheless, this strategy also induces uncertainty, and the selected similar ADs incur additional training costs without enough performance gains.
@@
To further address the aforementioned problems, we propose a novel detection method, named Adversarial Example Detection via Principal Adversarial Domain Adaptation (AED-PADA). As shown in Fig. 1(d), by selecting Principal Adversarial Domains (PADs) as the source domains, which significantly enlarges the coverage of the entire adversarial feature space and creates larger overlap with the target domain, we can offer superior generalization performance.
@@
Specifically, AED-PADA contains two stages, i.e., Principal Adversarial Domains Identification (PADI) and Principal Adversarial Domain Adaptation (PADA). In the stage of PADI, since the discrepancies between the adversarial examples from various adversarial attacks are quite different compared to these of the ordinary classification tasks, we exploit adversarial supervised contrastive learning(Adv-SCL) to construct distinguishable ADs. Then, the selection of the most representative ADs must meet two key criteria. Firstly, there should be a clear distinction between candidate ADs to avoid redundancy caused by selecting similar ADs. Secondly, the combination of the candidate ADs should cover as much of the entire feature space as possible. To select the most representative ADs, we propose a Coverage of Entire Feature Space (CEFS) metric. With our CEFS metric, the formed PADs possess broad coverage of the entire feature space, and thus effectively improve the likelihood of capturing the location of the unseen target AD(s).
In the stage of PADA, we pioneer to exploit the mechanism of Multi-source Unsupervised Domain Adaptation (MUDA) to effectively utilize the rich knowledge acquired by PADs, to detect the unseen adversarial examples in the target domain. The framework of PADA is compatible with various existing MUDA methods. Since typical MUDAs only focus on extracting the semantic features from the spatial domain, we propose an adversarial feature enhancement module to extract features from both the spatial and frequency domains to construct a more comprehensive representation of adversarial examples.
@@
  • We propose Principal Adversarial Domain Adaptation for detecting adversarial examples, by exploiting adversarial feature enhancement based Multi-source Unsupervised Domain Adaptation (MUDA), which is compatible with various existing MUDA methods. To the best of our knowledge, this is the first work to exploit MUDA for adversarial example detection.
2. Related Work
2.1. Adversarial Example Detection
@@
The majority of existing adversarial example detection methods rely on statistical features (HendrycksG17, ; magnet, ; FarAwayDis1, ; FarAwayDis2, ; Hidden4, ; Erase-and-Restore, ). They usually assume that the benign and adversarial examples originate from different distributions, and construct detectors based on the distinct statistical characteristics of these examples. Specifically, Grosse et al. (STest1, ) utilize Maximum Mean Discrepancy for the adversarial example detection. Li et al. employ Principal Component Analysis (PCA) to extract statistical feature, and construct a cascade classifier based on Support Vector Machines (SVMs) (PCAInconsistency, ). Feinman et al. (KDBU, ) carry out the detection based on Kernel Density (KD) and Bayesian-Uncertainty (BU) estimation. Ma et al. (lid, ) exploit the concept of Local Intrinsic Dimensionality (LID) to calculate the distance between the distribution of inputs and their neighbors. Lee et al. (md, ) utilize Guassian Discriminant Analysis (GDA) to model the difference between the benign and adversarial samples, and differentiate them based on Mahalanobis Distance (MD). Liu et al. (steg, ) point out that steganalysis could be applied to adversarial example detection, and propose a steganalysis-based detection method (Steg). Tian et al. (SID, ) reveal the inconsistency in the boundary fluctuations between the adversarial and benign examples, and construct Sensitivity Inconsistency Detector (SID) to identify the adversarial examples. Wang et al. (txt_advdetection, ) embed hidden-layer feature maps of DNNs into word vectors, and detect adversarial examples via Sentiment Analysis (SA).
Existing adversarial detection methods exhibit poor generalization because their training typically depends on a single known attack which vastly differs from unseen test attacks. In this paper, we propose a novel adversarial detection method, which can substantially increase the coverage of the entire adversarial feature space and create larger overlap with the test adversarial attacks.
2.2. Multi-source Unsupervised Domain Adaptation
Transfer learning (transfer_learning_survey1, ; transfer_learning_survey2, ; transfer_learning_survey3, ) is a deep learning technique which leverages knowledge acquired from the source task(s) to improve learning efficiency and performance on a related but different target task. Unsupervised domain adaptation (UDA) (unsuperviesed_domain_adaptation1, ) is a type of popular method in transfer learning which aims to migrate knowledge learned from the labeled source domain(s) to the target domain, where only unlabeled target data are available for training. Single-source Unsupervised Domain Adaptation (SUDA) (single-uda1, ; single-uda2, ; single-uda3, ) is widely explored in the previous research, which can transfer knowledge from one single source to one target domain. Compared to SUDA, Multi-source Unsupervised Domain Adaptation (MUDA) acquires richer information while introduces a new challenge, i.e., how to effectively bridge the domain gaps between all source domains and the target domain.
Various distribution alignment schemes have been proposed to achieve alignment between source and target domains. For example, Multiple Feature Spaces Adaptation Network (MFSAN) (aaai-mda, ) leverages Maximum Mean Discrepancy (MMD) to align the distributions of each pair of source and target domains in multiple specific feature spaces and aligns the outputs of classifiers by utilizing the domain-specific decision boundaries. Peng et al. (moment-mda, ) provide new theoretical insights specifically for moment matching to align the sources with each other and with the target. Owing to the development of generative adversarial networks, adversarial learning is widely used to find a domain-invariant feature space. It either focuses on approximating all combinations of pairwise domain discrepancies between each source and the target (cocktail, ; DARN, ) or uses a single domain discriminator (MIAN, ). Other explicit measures of discrepancy, such as Wasserstein distance (mdmn, ; Wasserstein, ), are also employed in MUDA to align the distribution of features. In addition to distribution alignment, the graph-matching metric (graph-muda1, ; graph-muda2, ) also considers the structural and geometric information, which achieves the alignment between the source and target domains by mapping both nodes and edges in a graph.
In this paper, we argue that the poor generalization performance of adversarial detection is due to the significant discrepancy between the source domains utilized for training and the target domain utilized for testing. Therefore, based on the MUDA approach, we propose a viable solution, Principal Adversarial Domain Adaptation, to reduce this great gap.
3. Methodology
@@
To improve the generalization ability of adversarial example detection, we propose a novel adversarial example detection method, named Adversarial Example Detection via Principal Adversarial Domain Adaptation (AED-PADA). The entire framework of our AED-PADA is shown in Fig. 2. AED-PADA contains two stages, Principal Adversarial Domains Identification (PADI) and Principal Adversarial Domain Adaptation (PADA). In the stage of PADI, we first incorporate adversarial supervised contrastive learning(Adv-SCL) to acquire distinguishable ADs. Then, we construct AD clustering to group ADs into different clusters. By proposing the Coverage of Entire Feature Space (CEFS) metric, we select the most representative ADs from each cluster to form PADs. In the stage of PADA, we propose an adversarial feature enhancement method based on the original MUDA method to effectively leverage PADs to detect the unseen adversarial attack methods. Note that Secs. 3.2, 3.3, 3.4 and 3.5 introduce PADI while Sec. 3.6 presents PADA.
3.1. Notations
@@
Suppose we have a labeled $C-$class classification dataset $\mathbb{D}=\{(x_{i},y_{i})\}_{i=1}^{N}$ with $N$ samples, where the label $y\in\{1,2,\ldots,C\}$. A classifier $f$ is trained on $\mathbb{D}$ to classify an input sample into one of the $C$ classes, $f(x)\rightarrow\mathbb{Z}_{C}$. Adversarial attack $\psi$ aims to fool $f$ into assigning incorrect labels via generating adversarial examples. We have a set of adversarial attack methods, $\Psi=\{\psi^{m}\}_{m=1}^{M}$, which consists of $M$ distinct adversarial attack methods. $\mathbb{D}^{m}=\{(x_{i}^{m},y_{i}^{m})\}_{i=1}^{N}$ is the adversarial dataset generated by the $m$-th adversarial attack method $\psi^{m}$, where $x_{i}^{m}$ denotes the $i$-th adversarial example generated by the $m$-th adversarial attack $\psi^{m}$, and $y_{i}^{m}$ presents its corresponding prediction label. We define $\varphi$ as, $\varphi(x_{i}^{m})=\psi^{m}$, which is a mapping from an adversarial example to its attack method. Consequently, we construct a set of adversarial examples $\mathcal{D}$ based on $\mathbb{D}$, which comprises $M$ types of adversarial examples, $\mathcal{D}=\{\mathbb{D}^{m}\}_{m=1}^{M}$.
Figure 2. Our AED-PADA framework contains two stages: (a) Principal Adversarial Domains Identification, which consists of Adversarial Domain Acquisition, Adversarial Domain Clustering and Principal Adversarial Domains Selection, and (b) Principal Adversarial Domain Adaptation.
3.2. Adversarial Domain Acquisition
@@
Typically, when we extract the features of the adversarial examples, which are generated from different untargeted attacks via common CNNs, the features tend to spread in an indistinguishable manner in the feature space. To acquire distinguishable representations of adversarial examples from different attacks, we exploit adversarial supervised contrastive learning (Adv-SCL) to extract features. Then, we form the Adversarial Domains (ADs), each of which is defined as the representations of all the adversarial examples generated from a particular adversarial attack.
@@
Based on the supervised contrastive learning (SCL) method (KhoslaTWSTIMLK20, ), Adv-SCL neglects the classification result of the adversarial examples and focuses sorely on identifying their generation methods. Specifically, each adversarial example $x_{i}^{m}$ in $\mathcal{D}$ is characterized by the method used to generate it, i.e., the adversarial attack method $\psi^{m}$. The $\psi^{m}$ of the adversarial example $x_{i}^{m}$ serves as a key discriminant for determining whether different $x_{i}^{m}$ are positive or negative samples. Here, a pair of examples from the same attack are considered as positive, while those from different attacks are considered as negative. This learning strategy amplifies the dissimilarities across examples from various attacks and generates a more appropriate representation.
@@
As shown in Fig. 2(a), the input of AD acquisition is the adversarial example set $\mathcal{D}$. Adv-SCL consists of an Encoder Network ($Enc$) and a Projection Network ($Proj$). $Enc(\cdot)$ extracts a feature vector from the input adversarial example $x_{i}^{m}$, and $Proj(\cdot)$ further projects this representation vector to an auxiliary vector $z_{i}=Proj(Enc(x_{i}^{m}))$, which will be discarded after training.
(transforming1, ; transform2, ) investigate that transformation strategies, such as cropping and rescaling, bit-depth reduction, JPEG compression, and randomization, have been used to defend against adversarial examples. Therefore, in order to prevent any potential ineffectiveness of adversarial samples caused by transformations, we only use normalization ($Norm$) as the data augmentation operation in the stage of AD Acquisition.
@@
We randomly sample $n$ example-label pairs in $\mathcal{D}$, $\{(x_{k}^{m_{k}},y_{k}^{m_{k}})\}_{k=1}^{n}$, where $m_{k}$ is the adversarial attack of the $k$-th adversarial example. The corresponding batch employed for training consists of $2n$ pairs, $\{(\tilde{x}_{l},\tilde{y}_{l})\}_{l=1}^{2n}$, where $\tilde{x}_{2k}$ and $\tilde{x}_{2k-1}$ are two views of $x_{k}^{m_{k}}$, and they are from the same adversarial attack method $\psi^{m_{k}}$. The loss function of Adv-SCL is defined as,
(1) $\mathcal{L}^{sup}=\sum_{i\in I}\mathcal{L}_{i}^{sup},$
(2) $\mathcal{L}_{i}^{sup}=\frac{-1}{|P(i)|}\sum_{p\in P(i)}\log\frac{\exp(z_{i}\cdot z_{p}/\tau)}{\sum_{a\in A(i)}\exp(z_{i}\cdot z_{a}/\tau)}.$
@@
Here, $i\in I=\{1,\ldots,2n\}$ denotes the index of the augmented samples. $A(i)$ is a subset of $I$ which includes all indexes except $i$. $P(i)=\{p\in A(i):\varphi(\tilde{x}_{p})=\varphi(\tilde{x}_{i})\}$ represents the indices of positive samples in the batch except $i$, and $|P(i)|$ stands for its cardinality. $\cdot$ denotes the dot product. $\tau$ is the temperature parameter which scales the similarity values.
3.3. Adversarial Domain Clustering